OpenStudy (anonymous):

Correct me if i'm right {n^2 e^-n} i think this result is diverge

OpenStudy (anonymous):

N0 \[ \frac {n^2}{e^n} \] Apply L'Hospital Rule to see that the sequence converges to zero.

OpenStudy (anonymous):

ok let me c

OpenStudy (anonymous):

if we use zero then is 0^2/e^0 = 0/1=0

OpenStudy (anonymous):

Apply it near \[ n=\infty\]

OpenStudy (anonymous):

you get \[ \frac {\infty} {\infty} \]

OpenStudy (anonymous):

lim of n-> infty

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k thx